OpenStudy (wilder.monday):

PLEASE HELP 1. ln x - ln(x-2)=5

OpenStudy (solomonzelman):

Some rules of logarithms: \(\color{#000000 }{ \displaystyle \ln(a)+\ln(b)=\ln\left(a \times b\right) }\) \(\color{#000000 }{ \displaystyle \ln(a)-\ln(b)=\ln\left(\frac{a}{b}\right) }\) \(\color{#000000 }{ \displaystyle a=a\times \ln(e)=\ln(e^a)}\) \(\color{#000000 }{ \displaystyle \ln(a)=b\quad \Longrightarrow \quad e^b=a }\)

OpenStudy (wilder.monday):

i'm still a little confused...

OpenStudy (solomonzelman):

If I had something like, (for example) \(\color{#000000 }{ \displaystyle \ln(z)-\ln(z+5)=10 }\) then, I would do this: \(\color{#000000 }{ \displaystyle \ln\left(\frac{z}{z+5}\right)=10 }\) (via the 2nd rule i posted) \(\color{#000000 }{ \displaystyle e^{10}=\frac{z}{z+5} }\) (via the last rule i posted) Then, algebra: (1) Multiply both sides times \(z+5\) \(\color{#000000 }{ \displaystyle (z+5)e^{10}=z }\) (2) expand the left side \(\color{#000000 }{ \displaystyle e^{10}z+5e^{10} =z }\) (3) subtract \(e^{10}z\) from both sides. \(\color{#000000 }{ \displaystyle 5e^{10} =z-e^{10}z }\) (4) factor the right side \(\color{#000000 }{ \displaystyle 5e^{10} =z }\) (5) divide by \(\color{#000000 }{ \displaystyle (1-e^{10}) }\) on both sides \(\color{#000000 }{ \displaystyle \frac{5e^{10} }{1-e^{10}}=z }\)

OpenStudy (solomonzelman):

Note that \(e\) is just a constant.

OpenStudy (wilder.monday):

oh ok thanks for clarifying i was confused with e part

OpenStudy (solomonzelman):

Also, very important property, \(\color{#000000 }{ \displaystyle \log_a(a)=1 }\) (this is true when "a" is any positive number, except 1) Same way, \(\color{#000000 }{ \displaystyle \ln(e)=\log_e(e)=1 }\) (because \(\color{#000000 }{ \displaystyle \ln }\) is just a log with base of "e")

OpenStudy (wilder.monday):

is e the base? in your explanation of how to solve the problem

OpenStudy (solomonzelman):

yes, everytime you see ln, it is same as log\(_e\), if that is what you are asking

OpenStudy (wilder.monday):

yes it is :)

OpenStudy (wilder.monday):

im doing the calculations now

OpenStudy (wilder.monday):

i got 2.0135....

OpenStudy (wilder.monday):

i checked it and it seems about right i think

OpenStudy (wilder.monday):

mind helping with some more?

OpenStudy (solomonzelman):

Yes, if I have time, but I need to ask you a question thougfh...

OpenStudy (wilder.monday):

yes?

OpenStudy (solomonzelman):

don't you need to find the exact solution for x? (not just an aproximation)

OpenStudy (wilder.monday):

no my teacher doesn't mind. i just put a couple numbers after the decimal point and she doesn't mind

OpenStudy (solomonzelman):

Ok:)

OpenStudy (wilder.monday):

i posted another question if you have time :)